# Cohomology of Fuchsian groups and Fourier interpolation

### v1

@@ 1. Introduction
The morphism $\Psi$ is in fact a homeomorphism of topological vector spaces with reference to natural topologies. We will give another proof of this theorem. The first step of this proof is to notice that the evaluation points $\sqrt{n}$ occur very naturally in the theory of the oscillator representation defined by Segal–Shale–Weil. Using this observation, the theorem can be reduced to computing the cohomology of a certain Fuchsian group with coefficients in this oscillator representation, and here we prove a more general statement:
@@
Let $G$ be $\mathrm{SL}_{2}(\mathbf{R})$ or a finite cover, $\Gamma$ a lattice in $G$, $W$ an irreducible infinite-dimensional $(\mathfrak{g},K)$-module, and $W^{*}_{-\infty}$ the distributional globalization of its dual (see 2.4). Then $H^{1}(\Gamma,W^{*}_{-\infty})$ is always finite dimensional, and in fact
@@
Some other interpolation consequences of Theorem 1.2 arise by replacing $\mathcal{S}$ by other spaces of functions carrying natural representations of $SL_{2}(\mathbf{R})$ and its finite covers: we discuss this in §6.4. For example any function $f\in C^{\infty}(\mathbf{R})$ for which $g(x):=f(1/x)$ extends to a smooth function at 00 is determined by $\hat{f}(n),\hat{g}(n)$ for nonzero $n\in\frac{1}{2}\mathbb{Z}$, together with the values and first derivatives of $f,g$ at 00.
@@
This follows from a Mayer-Vietoris type long exact sequence that computes the cohomology of the free group $\Gamma$, namely,
@@
  • The second proof is our original argument prior to learning of the work of Bunke-Olbrich just mentioned. It generalizes the standard way of computing $\Gamma$-cohomology with finite dimensional coefficients, as given in [BW00], to the infinite dimensional case – at least in cohomological degree $1$. Given the content of [BO98], we have permitted ourselves to abridge some tedious parts of our original argument, and reproduce here in detail the part that is perhaps most distinct from [BO98] – namely, we express the desired cohomology groups in terms of certain $\mathrm{Ext}$-groups of $(\mathfrak{g},K)$-modules and then compute these explicitly. Although we do not give full details here, this proof is quite explicit, and in particular it should be possible to produce an interpolation basis by explicating every step.
@@ 1.3. Questions
It is very interesting to ask about the situation where $\Gamma$ is not a lattice. Indeed, if one were to ask about an interpolation formula with evaluation points $1.1\sqrt{n}$, one is immediately led to similar questions for a discrete but infinite covolume subgroup of $\mathrm{SL}_{2}(\mathbf{R})$. Kulikov, Nazarov and Sodin [KNS23] have recently shown very general results about Fourier uniqueness that imply, in particular, that evaluating $f$ and $\hat{f}$ at $1.1\sqrt{n}$ do not suffice to determine $f$; it would be interesting to see if one can extract more precise information here by using the nonabelian harmonic analysis of the current paper.
@@
which is valid for general lattices $\Gamma$ in semisimple Lie groups $G$ and general irreducible (smooth, moderate growth) representations $V$ of $G$. Bunke and Olbrich have proved this in the cocompact case and our original argument proceeded by establishing the case $i=1$ for general lattices in $\mathrm{SL}_{2}(\mathbf{R})$.
@@ 1.4. Acknowledgements
We are very grateful to Matthew Emerton for his interest and input on the paper. These discussions led us to the work of Bunke and Olbrich. We also thank Henri Darmon and Joshua Mundinger for useful comments and suggestions.
M.G.G. was supported by the CRM and McGill University. A.V. was supported by an NSF-DMS grant during the preparation of this paper.
@@ 2.3. $(\mathfrak{g},K)$-modules
    If $V$ is irreducible, $\mathrm{Ext}^{1}_{(\mathfrak{g},K)}(W,V)$ is one-dimensional if $V\simeq V^{cl}$, and is zero otherwise.
@@ 3.1. Setup
One of the earliest versions of such a statement can be found in [GfGPS69, Chapter 1, §4]. For completeness we outline the proof, in our language, in Remark 4.3.
@@
The image of the map $\mathcal{C}-\lambda:\mathcal{A}_{K}\rightarrow\mathcal{A}_{K}$ is precisely the orthogonal complement to $\mathrm{Cusp}(\lambda)$ inside $\mathcal{A}_{K}$
@@ 5.1. Cusps
which describes a convex hyperbolic polygon which is (up to boundary) a fundamental domain for $\Gamma$ acting on $\mathbb{H}$; its pullback to $G$ via $g\mapsto g\cdot i$ is a fundamental domain for
$\Gamma\backslash G$, which will often be denoted by the same letter. In particular, $\mathfrak{F}$ can be decomposed in the following way, where the sets intersect only along their boundary:
@@ 6.1. The Weil representation
The vector $v_{1/2}:=e^{-\pi x^{2}}$ has $\kappa$-weight $1/2$ and Casimir eigenvalue $-3/8$ The other $K$-finite vectors in $\mathcal{S}$ are spanned by its Lie algebra translates; they have the form $q(x)e^{-\pi x^{2}}$ for $q$ an even polynomial, and have $\kappa$-weights $\frac{1}{2},\frac{5}{2},\frac{9}{2},\cdots$.
@@ 6.3. Conclusion of the proof
From 6.1, the $K$-finite vectors $\mathcal{S}_{K}$ are a realization of the $(\mathfrak{g},K)$-module of lowest weight $1/2$, whose complementary representation $(\mathcal{S}_{K})^{cl}$ is the $(\mathfrak{g},K)$-module of highest weight $-3/2$.
In general, a homomorphism from a lowest weight $(\mathfrak{g},K)$-module to any $(\mathfrak{g},K)$-module $W$ is uniquely specified by the image of the lowest weight vector, which can be an arbitrary element of $W$ killed by $m$; and the dual statement about highest weight modules is also valid.
@@
6.4. Variants: odd Schwartz functions, higher dimensions, $\mathbf{P}_{\mathbf{R}}^{1}$
@@ 6.4.1. Odd Schwartz functions
Indeed, the zeroth cohomology $H^{0}(\Gamma,\mathcal{T}^{*})$ is identified with the space of modular forms of weight $3/2$, a one-dimensional space spanned by $\theta^{3}$, as can be deduced from [CO76]. As for $H^{1}(\Gamma,\mathcal{T}^{*})$, its dimension is equal to the multiplicity of $\mathcal{T}^{cl}$ in the space of cusp forms on $\Gamma$. The representation $\mathcal{T}^{cl}$ has highest weight $-1/2$, and the vanishing of $H^{1}$ results from the absence of holomorphic cusp forms of weight $1/2$ on $\Gamma$ as in 6.2. We then deduce an interpolation theorem as in §1, noting that in addition to the $\delta_{n}$ the distributions $\phi\mapsto\phi^{\prime}(0)$ (resp. $\phi\mapsto\hat{\phi}^{\prime}(0)$) are also $e$- (resp. $f$-)invariant.
Arguing as in §1.1 recovers a non-explicit version of the interpolation theorem of Radchenko-Viazovska for odd Schwartz functions, see [RV19, Theorem 7].
6.4.2. Schwartz functions on $\mathbf{R}^{d}$
@@
6.4.3. Smooth functions on the projective real line
We consider the space $C^{\infty}(\mathbf{P}^{1}_{\mathbf{R}})$ of smooth functions on $\mathbf{P}^{1}_{\mathbf{R}}=\mathbf{R}\cup\infty$, which we may think of equivalently as either $\mathbf{R}^{\times}$-invariant smooth functions $\Phi(x,y)$ on $\mathbf{R}^{2}-\{0\}$ or smooth functions $h(x)$ on $\mathbf{R}$ with an asymptotic expansion $h\sim c_{0}+c_{1}/x+\dots$ as $x\rightarrow\infty$; to pass from the former to the latter we use $h(x)=\Phi(x,1)$. We define for $\Phi\in C^{\infty}(\mathbf{P}^{1}_{\mathbf{R}})$ :
@@ 6.4.3. Smooth functions on the projective real line
Theorem 6.3.
The map $h\mapsto(a_{n},b_{n},h(0),h^{\prime}(0),h(\infty),h^{\prime}(\infty))$ defines topological isomorphisms of $C^{\infty}(\mathbf{P}^{1}_{\mathbf{R}})$ with a codimension $5$ subspace of $\mathfrak{s}^{2}\oplus\mathbf{C}^{4}$.
In principle, the $5$ constraints can be explicated by means of the study of the boundary values of weight $2$ Eisenstein series, but the details are a little delicate, and we did not carry out the computation.
Proof.
@@
We obtain Theorem 6.3 in a similar way to Theorem 1.1 – namely, by applying Theorem 1.2 for the same $\Gamma$, but with a different coefficient system. Namely, we consider the space $W=C^{\infty}(\mathbf{P}^{1}_{\mathbf{R}})$ and apply Theorem 1.2 to the $(\mathfrak{g},K)$-module $W_{K}$.
Here, the $e$-invariants on $W^{*}$ are spanned by the $\Phi\mapsto a_{n}(\Phi)$ together with $\Phi(1,0)$ and $\partial_{y}\Phi(1,0)$ (i.e., $h(\infty)$ and $h^{\prime}(\infty)$). To check this, take an arbitrary $e$-invariant distribution $\mathcal{D}$ on $W$; its restriction to $C^{\infty}_{c}(\mathbf{R})$ is a periodic distribution under $x\mapsto x+2$ and is thus readily verified to be a linear combination of the distributions $a_{n}$ for $n\in\mathbf{Z}$. We may therefore suppose that $\mathcal{D}$ vanishes on $C^{\infty}_{c}(\mathbf{R})$, and so is supported at $\infty$, but it is then a linear combination of the various coefficients $c_{i}$ in the asymptotic expansion of
$h(x)\sim c_{0}+\frac{c_{-1}}{x}+\frac{c_{-2}}{x^{2}}+\dots$
and a computation shows that only $c_{0},c_{-1}$ are invariant under the translation map $x\mapsto x+2$. Symmetrically, $(W^{*})^{f}$ is spanned by the functionals $b_{n}(\Phi),h(0),h^{\prime}(0)$. Therefore, to prove the theorem, it is enough to show that $(W^{*})^{e}$ and $(W^{*})^{f}$ add up to $W^{*}$, and intersect in a $5$-dimensional space.
The space $W$ is identified with a reducible principal series of $SL_{2}(\mathbf{R})$, namely with the extension $\mathbf{C}\to W\to D_{2}^{\pm}$ of the discrete series of weight ±$\pm 2$ by the trivial representation. Either via Theorem 1.2 (plus an auxiliary computation) or directly from the results of [BO98, §8] one computes that
(54) $\mathrm{dim}H^{1}(\Gamma,W^{*})=2,\qquad H^{0}(\Gamma,W^{*})=5.$
in fact the natural map $W^{*}\rightarrow\mathbf{C}$ induces an isomorphism on $H^{1}$.
We apply again the Mayer-Vietoris sequence (5). It becomes, using subscripts and superscripts to denote coinvariants and invariants respectively,
(55) $0\to H^{0}(\Gamma,W^{*})\to(W^{*})^{e}\oplus(W^{*})^{f}\to(W^{*})\to H^{1}(\Gamma,W^{*})\to W^{*}_{e}\oplus W^{*}_{f}\to 0$
The constant function in $W$ is $e$-invariant and $f$-invariant. By evaluating at it we see that both $W^{*}_{e}$ and $W^{*}_{f}$ are at least one-dimensional. But then comparing (55) and (54) we find that the surjective map $H^{1}(\Gamma,W^{*})\to W^{*}_{e}\oplus W^{*}_{f}$ must be an isomorphism, and so $(W^{*})^{e}\oplus(W^{*})^{f}\to(W^{*})$ is surjective with $5$-dimensional kernel as required. ∎

### v2

@@ 1. Introduction
This is an abstract interpolation result: The statement implies the existence of a universal formula that computes any value $\phi(x)$ of any even Schwartz function $\phi$ as a linear combination $\sum a_{n}(x)\phi(\sqrt{n})+\sum\hat{a}_{n}(x)\hat{\phi}(\sqrt{n})$ for some $a_{n}(x),\hat{a}_{n}(x)$, but does not specify what those functions are. By contrast, Radchenko and Viazovska first write down this explicit interpolation formula, and then deduce Theorem 1.1 from it. In a sense, what is accomplished in the present paper is to separate the abstract content of this interpolation result from its computational aspect.
The morphism $\Psi$ is in fact a homeomorphism of topological vector spaces with reference to natural topologies. We will give another proof of this theorem. The first step of this proof is to notice that the evaluation points $\sqrt{n}$ occur very naturally in the theory of the oscillator representation defined by Segal–Shale–Weil (see [Cha12] or [LV13] for introductions). Using this observation, the theorem can be reduced to computing the cohomology of a certain Fuchsian group with coefficients in this oscillator representation, and here we prove a more general statement:
@@
Let $G$ be $\mathrm{SL}_{2}(\mathbf{R})$ or a finite cover thereof, $\Gamma$ a lattice in $G$, $W$ an irreducible infinite-dimensional $(\mathfrak{g},K)$-module, and $W^{*}_{-\infty}$ the distributional globalization of its dual (see 2.4). Then $H^{1}(\Gamma,W^{*}_{-\infty})$ is always finite dimensional, and in fact
@@
Some other interpolation consequences of Theorem 1.2, where interpolation is understood in the abstract sense as discussed after Theorem 1.1, arise by replacing $\mathcal{S}$ by other spaces of functions carrying natural representations of $SL_{2}(\mathbf{R})$ and its finite covers: we discuss this in §6.4. For example, Hedenmalm and Montes-Rodríguez [HMR11] have shown that the functions $e^{i\pi\alpha nt}$, $e^{i\pi\beta n/t}$ are weakly dense in $L^{\infty}$ if and only if $\alpha\beta=1$. We will show that an interpolation result holds at the transition point $\alpha\beta=1$; we thank the referee for bringing [HMR11] to our attention.
@@
This follows from a Mayer-Vietoris type long exact sequence that computes the cohomology of the free group $\Gamma$ [Bro12, II-III], namely,
@@
  • The second proof is our original argument prior to learning of the work of Bunke-Olbrich just mentioned. It generalizes the standard way of computing $\Gamma$-cohomology with finite dimensional coefficients, as given in [BW00], to the infinite dimensional case – at least in cohomological degree $1$. Given the content of [BO98], we have permitted ourselves to abridge some tedious parts of our original argument, and reproduce here in detail the part that is perhaps most distinct from [BO98] – namely, we express the desired cohomology groups in terms of certain $\mathrm{Ext}$-groups of $(\mathfrak{g},K)$-modules and then compute these explicitly.
@@ 1.3. Questions
As we have noted, we prove an abstract interpolation result. Can one recover the explicit formula for the interpolating functions, as given in [RV19], from this approach? It seems to us that our proof is sufficiently explicit that this is, at least, plausible.
It is very interesting to ask about the situation where $\Gamma$ is not a lattice. Indeed, if one were to ask about an interpolation formula with evaluation points $0.9\sqrt{n}$, one is immediately led to similar questions for a discrete but infinite covolume subgroup of $\mathrm{SL}_{2}(\mathbf{R})$, whereas considering $1.1\sqrt{n}$ leads to considering a non-discrete subgroup. Note that Kulikov, Nazarov and Sodin [KNS23] have recently shown very general results about Fourier uniqueness that imply, in particular, that evaluating $f$ and $\hat{f}$ at $1.1\sqrt{n}$ do not suffice to determine $f$, but that evaluating them at $0.9\sqrt{n}$ does.
@@
which is valid for general lattices $\Gamma$ in semisimple Lie groups $G$ and general irreducible (smooth, moderate growth) representations $V$ of $G$. Bunke and Olbrich have proved this in the cocompact case, and our original argument proceeded by establishing the case $i=1$ for general lattices in $\mathrm{SL}_{2}(\mathbf{R})$. Also, Deitmar and Hilgert prove [DH05, Corollary 3.3] a result of this type in great generality, but with the space of automorphic forms replaced by the larger space $C^{\infty}(\Gamma\backslash G)$ without growth constraints.
@@ 1.4. Acknowledgements
We are very grateful to Matthew Emerton for his interest and input on the paper. These discussions led us to the work of Bunke and Olbrich. We also thank Henri Darmon and Joshua Mundinger for useful comments and suggestions. We thank the referees for comments and suggestions that led to improvements in the exposition.
M.G-G. was supported by the CRM and McGill University. A.V. was supported by an NSF-DMS grant during the preparation of this paper.
@@ 2.3. $(\mathfrak{g},K)$-modules
    If $V$ is irreducible, $\mathrm{Ext}^{1}_{(\mathfrak{g},K)}(W,V)$ is one-dimensional if $V\simeq W^{cl}$, and is zero otherwise.
@@ 3.1. Setup
One of the earliest versions of such a statement can be found in [GGPS69, Chapter 1, §4]. For completeness we outline the proof, in our language, in Remark 4.3.
@@
The image of the map $\mathcal{C}-\lambda:\mathcal{A}_{K}\rightarrow\mathcal{A}_{K}$ is precisely the orthogonal complement to $\mathrm{Cusp}(\lambda)$ inside $\mathcal{A}_{K}$.
@@ 5.1. Cusps
which describes a convex hyperbolic polygon which is (up to boundary) a fundamental domain for $\Gamma$ acting on $\mathbb{H}$; its pullback to $G$ via $g\mapsto g\cdot i$ is a fundamental domain for $\Gamma\backslash G$, which will often be denoted by the same letter. In particular, $\mathfrak{F}$ can be decomposed in the following way, where the sets intersect only along their boundary:
@@ 6.1. The Weil representation
The vector $v_{1/2}:=e^{-\pi x^{2}}$ has $\kappa$-weight $1/2$ and Casimir eigenvalue $-3/8$. The other $K$-finite vectors in $\mathcal{S}$ are spanned by its Lie algebra translates; they have the form $q(x)e^{-\pi x^{2}}$ for $q$ an even polynomial, and have $\kappa$-weights $\frac{1}{2},\frac{5}{2},\frac{9}{2},\cdots$.
@@ 6.3. Conclusion of the proof
From 6.1, the $K$-finite vectors $\mathcal{S}_{K}$ are a realization of the $(\mathfrak{g},K)$-module of lowest weight $1/2$, whose complementary representation $(\mathcal{S}_{K})^{cl}$ is the $(\mathfrak{g},K)$-module of highest weight $-3/2$. In general, a homomorphism from a lowest weight $(\mathfrak{g},K)$-module to any $(\mathfrak{g},K)$-module $W$ is uniquely specified by the image of the lowest weight vector, which can be an arbitrary element of $W$ killed by $m$; and the dual statement about highest weight modules is also valid.
@@
6.4. Variants: odd Schwartz functions, higher dimensions, Heisenberg uniqueness.
@@ 6.4.1. Odd Schwartz functions
Indeed, the zeroth cohomology $H^{0}(\Gamma,\mathcal{T}^{*})$ is identified with the space of modular forms of weight $3/2$, a one-dimensional space spanned by $\theta^{3}$, as can be deduced from [CO76]. As for $H^{1}(\Gamma,\mathcal{T}^{*})$, its dimension is equal to the multiplicity of $\mathcal{T}^{cl}$ in the space of cusp forms on $\Gamma$. The representation $\mathcal{T}^{cl}$ has highest weight $-1/2$, and the vanishing of $H^{1}$ results from the absence of holomorphic cusp forms of weight $1/2$ on $\Gamma$ as in 6.2. We then deduce an interpolation theorem as in §1, noting that in addition to the $\delta_{n}$ the distributions $\phi\mapsto\phi^{\prime}(0)$ (resp. $\phi\mapsto\hat{\phi}^{\prime}(0)$) are also $e$- (resp. $f$-)invariant. Arguing as in §1.1 recovers a non-explicit version of the interpolation theorem of Radchenko-Viazovska for odd Schwartz functions, see [RV19, Thm. 7].
6.4.2. Radial Schwartz functions on $\mathbf{R}^{d}$
@@
6.4.3. Heisenberg uniqueness
A result of Hedenmalm and Montes-Rodríguez [HMR11] asserts that the map
@@ 6.4.3. Heisenberg uniqueness
is injective if and only if $\alpha\beta\leq 1$. In their terminology, this yields an example of a “Heisenberg uniqueness pair.” We thank the referee for bringing this result to our attention. Using our techniques, we show that an abstract interpolation formula – admittedly, on a eccentric function space – holds at the transition point $\alpha\beta=1$.
Theorem 6.3.
@@
$h\mapsto((a_{n}),(b_{n}),h(0),\lim_{x\rightarrow\infty}x^{2}h(x))$
defines a linear isomorphism of $\mathcal{H}$ with a codimension $3$ subspace $S$ of [^6] Note that integration by parts shows that $a_{n},b_{n}$ indeed belong to the space $\mathfrak{s}$ of sequences with rapid decay, introduced in §1. $\mathfrak{s}^{2}\oplus\mathbf{C}^{2}$.
In this form, this neither implies nor is implied by the results of [HMR11], but it would be interesting to see if our methods can give results closer to theirs, e.g. by considering different completions of the underlying representation.
@@
We obtain Theorem 6.3 in a similar way to Theorem 1.1 – namely, by applying Theorem 1.2 for the same $\Gamma$, but with a different coefficient system. Note that we can and will assume that $\alpha=\beta=1$ by rescaling. We now consider the space $W$ of smooth $1$-forms on $\mathbf{P}^{1}_{\mathbf{R}}=\mathbf{R}\cup\infty$, which we may think of equivalently as smooth functions $\Phi(x,y)$ on $\mathbf{R}^{2}-\{0\}$ satisfying
$\Phi(\lambda x,\lambda y)=\lambda^{-2}\Phi(x,y).$
The $1$-form on $\mathbf{P}^{1}_{\mathbf{R}}$ associated to $\Phi$ is characterized by the fact that, when pulled back to $\mathbf{R}^{2}-\{0\}$, it gives the $1$-form $\Phi(x,y)(xdy-ydx)$. Write
[…]
Write $h(x)=\Phi(x,1)$. We note that that $x^{-2}h(1/x)=\Phi(1,x)$, and so extends over 00. The map $\Phi\mapsto h(x)=\Phi(x,1)$ thus identifies $W$ with the space $\mathcal{H}$ described in the theorem. We are reduced then to proving:
Claim: the rule
(56) $\Phi\mapsto(a_{n},b_{n},\Phi(0,1),\Phi(1,0))$
defines an isomorphism of $W$ with a codimension $3$ subspace of $\mathfrak{s}^{2}\oplus\mathbf{C}^{2}$.
Proof of Claim: We apply Theorem 1.2 to the $(\mathfrak{g},K)$-module $W_{K}$; the distribution globalization “$W^{*}_{-\infty}$” that appears in Theorem 1.2 is simply the topological dual $W^{*}$ to $W$.
To analyze the $e$-invariants on $W^{*}$, take an arbitrary $e$-invariant distribution $\mathcal{D}$ on $W$. The identification $\Phi\mapsto h$ between the space of $-2$-homogeneous $\Phi$ and $h\in\mathcal{H}$ contains $C^{\infty}_{c}(\mathbf{R})$ in its image; thus, we can consider $\mathcal{D}$ as a distribution on the real line, i.e., given any $h\in C^{\infty}_{c}(\mathbf{R})$, we form the corresponding $\Phi$ and evaluate $\mathcal{D}$ on it. The result is a periodic distribution under $x\mapsto x+2$ which must be in the closed subspace spanned by the $a_{n}$ for $n\in\mathbf{Z}$ – write this distribution $\sum c(n)a_{n}$. Then the difference $\mathcal{D}-\sum c(n)a_{n}$ vanishes on $C^{\infty}_{c}(\mathbf{R})$, and is therefore a linear combination of the Taylor coefficients of $\Phi(1,y)$ at $y=0$; the only such distribution that is invariant under $e$ is $\Phi\mapsto\Phi(1,0)$. It follows that $(W^{*})^{e}$ is spanned topologically by the $a_{n}$ and evaluation at $(1,0)$. Similarly, $(W^{*})^{f}$ is spanned topologically by the $b_{n}$ and evaluation at $(0,1)$.
We will now compute the cohomology of $\Gamma$ on $W^{*}$.
The space $W$ is identified with a reducible principal series of $SL_{2}(\mathbf{R})$ which is an extension $D_{2}^{+}\oplus D_{2}^{-}\to W\to\mathbf{C}$, where $D_{2}^{\pm}$ are the holomorphic and antiholomorphic discrete series of weight $2$; the map $W\rightarrow\mathbf{C}$ is the integration over $\mathbf{P}^{1}_{\mathbf{R}}$. Now Theorem 1.2 implies that $H^{1}(\Gamma,(D_{2}^{\pm})^{*})$ vanishes, whereas $H^{0}(\Gamma,(D_{2}^{\pm})^{*})$ has dimension $2$ in both the $+$ and -- case. There is therefore an exact sequence
$0\rightarrow\mathbf{C}\rightarrow H^{0}(\Gamma,W^{*})\rightarrow\mathbf{C}^{4}\rightarrow H^{1}(\Gamma,\mathbf{C})\rightarrow H^{1}(\Gamma,W^{*})\rightarrow 0$
The map $\mathbf{C}^{4}\rightarrow H^{1}(\Gamma,\mathbf{C})$ is surjective with two-dimensional kernel, for it amounts to the the map from the $4$-dimensional space of (holomorphic and antiholomorphic) Eisenstein series for $\Gamma$ to the two-dimensional cohomology. 777 Indeed, this map records the obstruction to extending an embedding of $D_{2}^{+}\oplus D_{2}^{-}$ into the space of automorphic forms, to the larger space $W$. An embedding of $D_{2}^{+}\oplus D_{2}^{-}$ into the space of automorphic forms corresponds to a pair $(f,g)$ of a holomorphic and antiholomorphic $1$-form, and it extends to $W$ when $fdz+g\overline{dz}$ is an exact differential. This proves that $H^{1}(\Gamma,W^{*})$ vanishes, whereas $H^{0}(\Gamma,W^{*})$ is $3$-dimensional.
We now apply the Mayer-Vietoris sequence (5). In our current context, it implies that $(W^{*})^{e}$ and $(W^{*})^{f}$ span all of $W^{*}$, and their intersection is precisely $3$-dimensional. This concludes the proof of the claim. ∎
It may be of interest to describe the three linear constraints that define this codimension $3$ subspace. We follow the notations above. The invariants of $\Gamma$ on $W^{*}$ have, as basis $A,I,J$ where
$A(\Phi)=\int_{\mathbf{P}^{1}}\Phi$
$I(\Phi)=\sum_{(m,n)\neq(0,0)}\Phi(m,n)-2\sum_{2|n}\Phi(m,n)$
$J(\Phi)=\sum_{(m,n)\neq(0,0)}\Phi(m,n)-2\sum_{2|m}\Phi(m,n)$
where in both cases the sum is conditionally convergent (e.g., one can sum over large discs of increasing radii). Then $A$ corresponds to the relation $a_{0}=b_{0}$, whereas both $I$ and $J$ give rise to a relation by expanding the stated intertwiner in two different ways. For example, we compute $I(\Phi)$ in two ways, firstly by summing first over $n$:
$\sum_{m\neq 0}m^{-2}\sum_{n}(\Phi(1,\frac{n}{m})-2\Phi(1,\frac{2n}{m}))+2(\sum n^{-2}-2\sum(2n)^{-2})\Phi(0,1)$
=$\stackrel{{\scriptstyle\textrm{P.S.}}}{{=}}-\sum_{m\neq 0,t\in\mathbf{Z}}|m|^{-1}b_{m(2t+1)}+\frac{\pi^{2}}{6}\Phi(0,1)$
where P.S. stands for Poisson summation, and secondly by summing first over $m$:
$\sum_{n\neq 0}\left(n^{-2}\sum_{m}\Phi(\frac{m}{n},1)-\sum_{m}2(2n)^{-2}\Phi(\frac{m}{2n},1)\right)2(\sum m^{-2}-2\sum(m)^{-2})\Phi(1,0)$
=$\stackrel{{\scriptstyle\textrm{P.S.}}}{{=}}\sum|n|^{-1}a_{n(4t+2)}-\frac{\pi^{2}}{3}\Phi(1,0).$
Thus we find that the image of $W$ is cut out by the three relations $a_{0}=b_{0}$,
$\frac{\pi^{2}}{3}\Phi(1,0)+\frac{\pi^{2}}{6}\Phi(0,1)=\sum_{m\neq 0,t\in\mathbf{Z}}|m|^{-1}b_{m(2t+1)}+\sum_{n\neq 0,t\in\mathbf{Z}}|n|^{-1}a_{n(4t+2)},$
and dually
$\frac{\pi^{2}}{6}\Phi(1,0)+\frac{\pi^{2}}{3}\Phi(0,1)=\sum_{m\neq 0,t\in\mathbf{Z}}|m|^{-1}a_{m(2t+1)}+\sum_{n\neq 0,t\in\mathbf{Z}}|n|^{-1}b_{n(4t+2)}.$